### v1

# Thompson sampling MAB multi-armed bandit

Abstract
This paper focuses on selecting the arm with the highest variance from a set of $K$ independent arms. Specifically, we focus on two settings: (i) regret setting, that penalizes the number of pulls of suboptimal arms in terms of variance, and (ii) fixed-budget best arm identification (BAI) setting, that evaluates the ability of an algorithm to determine the arm with the highest variance after a fixed number of pulls. We develop a novel online algorithm called UCB-VV for the regret setting and show that its upper bound on regret for bounded rewards evolves as $\mathcal{O}\left(\log{n}\right)$ where $n$ is the horizon. By deriving the lower bound on the regret, we show that UCB-VV is order optimal. For the fixed budget BAI setting and propose the SHVV algorithm. We show that the upper bound of the error probability of SHVV evolves as $\exp\left(-\frac{n}{\log(K)H}\right)$, where $H$ represents the complexity of the problem, and this rate matches the corresponding lower bound. We extend the framework from bounded distributions to sub-Gaussian distributions using a novel concentration inequality on the sample variance. Leveraging the same, we derive a concentration inequality for the empirical Sharpe ratio (SR) for sub-Gaussian distributions, which was previously unknown in the literature. Empirical simulations show that UCB-VV consistently outperforms $\epsilon$-greedy across different sub-optimality gaps though it is surpassed by VTS, which exhibits the lowest regret, albeit lacking in theoretical guarantees. We also illustrate the superior performance of SHVV, for a fixed budget setting under 6 different setups against uniform sampling. Finally, we conduct a case study to empirically evaluate the performance of the UCB-VV and SHVV in call option trading on $100$ stocks generated using geometric Brownian motion (GBM).
Index Terms:
Multi-Armed bandits, risk, variance, regret, best arm identification.
I Introduction and Motivation
Historically, variance has served as the first and foundational risk measure in portfolio optimization. The modern portfolio theory is based upon the classical Markowitz model, which also employs variance as a measure of risk [1, 2, 3]. In economic models, the decision-makers account for uncertainty and variance to maintain robustness [4]. In decision-making and optimization, identifying and mitigating sources of high variance is crucial to ensuring stability and robustness in outcomes [5, 6, 7]. Variance also plays a critical role in strategies involving trading call options, as higher variance increases the likelihood of the underlying asset’s price exceeding the strike price, thereby enhancing the intrinsic value of the option [8]. Also, it is a well-known fact that a call option on a volatile stock is more valuable than one that is less volatile [9].
Recent advances in risk-sensitive decision making involve more sophisticated risk measures such as value-at-risk (VaR) and conditional value-at-risk (CVaR) [10, 11, 12, 13, 14]. Unlike VaR and CVaR, which rely on tail-specific assumptions and involve complex computations related to probability thresholds and conditional expectations, variance is distribution-agnostic and computationally lightweight. From the perspective of selecting the arm with the highest variance, we study two settings of the multi-armed bandit (MAB) framework - regret minimization and BAI. Specifically, the regret setting attempts to minimize the number of pulls of sub-optimal arms in terms of the variance. On the contrary, the BAI setting attempts to determine the arm with the highest variance with a high probability.
For portfolio managers, maximizing the frequency of selecting the highest variance option unlocks opportunities to optimize returns while leveraging the asymmetrical payoff structure of call options from favorable price movements. For example, for a call option buyer, a higher variance is particularly advantageous as it increases the potential for substantial gains while capping the maximum loss to the premium paid for the option. In this paper, we investigate this aspect under the regret setting by developing a bandit algorithm UCB-VV for call option trading using the variance associated with a stock. It must be noted that the results of our framework can be extended to the setting to pick the arm with the minimum variance as well.
Unlike the regret setting, the fixed-budget BAI focuses on identifying the best arm with minimum error for a fixed number of samples [15, 16, 17]. This is particularly relevant in scenarios like A/B testing, clinical trials, or recommendation systems, where the ultimate goal is to find the single best option. In finance, high-variance assets introduce instability and expose portfolios to erratic swings, requiring their identification and possible exclusion.
I-A Related work
I-A1 Literature in regret and BAI setting
In the regret setting, [18] investigates UCB-V, a variant of the upper confidence bound (UCB) algorithm for stochastic MAB that incorporates empirical variance estimates of arm payoffs to pick the arm with the highest mean. The algorithm performs better than traditional methods, especially when suboptimal arms have low variance. However, it depends on an unavoidable upper-limit $u$, which can reduce efficiency when $u$ is large but actual variances are small. The authors in [19] present E-UCBV, a variance-aware UCB algorithm that achieves tighter regret bounds than UCB1. By incorporating efficient arm elimination, E-UCBV demonstrates superior performance in scenarios with small gaps and low-variance suboptimal arms. However, the algorithm has notable limitations, including high complexity in regret bounds and constraints on $n$. The authors in [20, 21] address risk-aware decision-making and design algorithms for mean-variance MAB.
The BAI problem has been extensively investigated across a broad range of settings, e.g, for linear bandits [22], cost-aware BAI [23], as well as risk-averse bandits [24]. In risk-averse BAI problem, [25] considers VaR-optimization. Also, [24], [26], optimizes a mean-CVaR framework within a fixed budget framework. They propose distribution oblivious algorithms, i.e, they do not require prior knowledge of the arms’ underlying distributions and are asymptotically near-optimal. However, these algorithms come with weaker performance guarantees compared to those that leverage prior knowledge of the arm distributions, highlighting an inherent trade-off between robustness and efficiency. Variance optimization, in particular, often necessitates stronger assumptions about the reward distributions (e.g., sub-Gaussianity) to ensure accurate and reliable estimation. Without such assumptions, variance estimates can become highly sensitive to noise and unreliable.
I-A2 Literature in finance
In the context of finance, variance is used to model risk premium, providing insights into market sentiment and investor behavior. Specifically, the variance risk premium (a compensation that investors demand for bearing the risk associated with the volatility in the returns of an asset) plays a critical role in both theoretical and practical applications. This concept has been leveraged to develop trading strategies, such as selling volatility when the premium is high, as demonstrated in [27, 28]. Beyond its application in trading strategies, variance is also instrumental in understanding cross-market dynamics. For instance, Hamao et al. [29] investigate how volatility in one market, such as Japan, influences volatility in another, such as the U.S., using variance as a metric to assess volatility spillovers. Similarly, Engle et. al [30] examine intraday volatility in foreign exchange markets, employing variance to model the clustering of volatility over time. This clustering, a manifestation of the heteroscedastic behavior, is effectively captured by variance. Variance has been central to research on asset pricing and portfolio management. French et al. [31] explore the relationship between expected stock returns and volatility, with volatility quantified by variance, thereby informing optimal portfolio construction and asset pricing. Collectively, these studies illustrate the multifaceted role of variance in financial research and practice.
I-B Contributions and Organization
  1. 1.
    Regret algorithm and optimality: First, for the regret setting, we propose a UCB-type algorithm – UCB-VV that maximizes the frequency of selecting the arm with the highest variance. We show that the upper bound on the regret for UCB-VV is $\mathcal{O}(\log{n})$, and derive the corresponding information-theoretic lower bound to demonstrate that UCB-VV is order optimal. This is presented in Section II.
  2. 2.
    BAI algorithm and optimality: Next, we propose an arm-elimination based algorithm called SHVV to identify the arm with the highest variance within a fixed budget of pulls. Similar to the regret case, we derive the information-theoretic lower bound on the probability of error for the highest variance BAI and show that SHVV is order optimal. This is presented in Section III.
    Key novelty in 1) and 2): This is the first work that develops algorithms for selecting the arm with the highest variance and proves their order optimality.
  3. 3.
    Extension to sub-Gaussian distributions and applications to Sharpe ratio (SR): While the above results consider bounded distributions, we prove that our framework is more general and can be applied to sub-Gaussian distributions. For this, we derive a novel Bernstein-type concentration inequality for the sample variance of a sub-Gaussian distribution. Leveraging the same, we derive a novel concentration bound for the sample SR. We demonstrate the application of this result to devise a UCB-type algorithm to select the arm with the highest SR and provide a sketch for bounding its regret. This discussion is presented in Section IV.
    Key novelty in 3): To the best of our knowledge, this is the first work that provides a path-independent concentration bound (unlike [32]) for the sample SR for sub-Gaussian distributions. The resulting algorithm of SR optimality and its regret bound has not been reported in the literature.
  4. 4.
    Practical Application: Finally, following the theoretical development, in Section V, we compare the performance of UCB-VV with algorithms like $\epsilon$-greedy, VTS, and KL-UCB. We also demonstrate the superior performance of SHVV for a fixed budget setting under 6 different setups. Then, in Section VI, we validate the efficacy of the proposed algorithms (both UCB-VV and SHVV) on a real-world application – call option trading of GBM simulated $100$ stocks.
II Problem formulation for regret
Consider a MAB framework with $K\geq 2$ arms, a single player, and a time horizon $n$. We assume that the underlying rewards distributions of the arms are $\nu_{1},\nu_{2},\dots,\nu_{K}$, each of which is bounded [^1] In Section IV we relax the boundedness assumption and demonstrate the extension of our framework to sub-Gaussian distributions. in $[l,u]$. The means, $\{\mu_{i}\}$, and the variances, $\{\sigma_{i}^{2}\}$, of the distributions are fixed but unknown to the player. When playing an arm $i$ at time $t$, a random reward $X_{i}(t)\sim\nu_{i}$. The reward samples are independent across arms and time slots. A policy $\pi$ orchestrates the action of the player at each time instant and is driven by the historical data. Let $s_{i}(n)$ represent the number of plays of an arm $i$ until (and including) time $n$ by the player. In a standard stochastic bandit problem, the regret setting focuses on optimizing the trade-off between exploration and exploitation to maximize the expected cumulative reward over time. Without loss of generality, we assume that the optimal arm is unique and $\sigma^{2}_{\ast}=\max_{i\in[K]}\sigma_{i}^{2}$.
Definition 1.
The biased empirical variance of an arm $i$ at time $t$, calculated from the i.i.d. samples $\{X_{i}(s_{i}(q))\}_{q=1}^{t}$ drawn from its distribution $\nu_{i}$ given as
$\bar{V}_{i}(t)=\frac{1}{s_{i}(t)}\sum_{q=1}^{s_{i}(t)}\left(X_{i}(s_{i}(q))-\bar{\mu}_{i}(s_{i}(t))\right)^{2},$ (1)
where $\bar{\mu}_{i}(s_{i}(t))=\frac{1}{t}\sum_{q=1}^{t}X_{i}(s_{i}(q)).$
Then, for a policy $\pi$ and horizon $n$, we define the regret as
$\mathbb{E}[R(n)]=\sum_{i=1}^{K}\mathbb{E}[s_{i}(n)]\delta_{i},$ (2)
where $\delta_{i}=\sigma^{2}_{*}-\sigma_{i}^{2}$ denotes the sub-optimality gap of an arm $i$ with respect to the optimal arm.
Remark 1.
The regret defined in (2) deviates from the classical definition of regret (e.g., see [33]) that considers the regret as the expected deviation of the metric obtained by the algorithm from its optimal value. Unlike the mean, for maximizing the variance, the optimal policy may not be equivalent to sampling the arm with the highest variance at all times, but consists of strategic switching between the arms [34]. However, since our focus is on selection of the arm with the highest variance, and penalizing a policy each time it selects a variance sub-optimal arm, (2) is an appropriate choice of measuring the performance.
Note that employing a biased estimate of variance does not have any impact on our regret guarantees. As $t$ increases, the distinction between $t-1$ and $t$ becomes increasingly negligible due to their asymptotic equivalence. This adjustment impacts all arms equally without altering their relative ordering.
II-A UCB-VV Algorithm
We introduce UCB-VV (presented in Algorithm 1) to pull an arm with the highest variance most frequently. The algorithm pulls the arm with the highest index $B_{i}(t)_{\rm VV}$, where
$B_{i}(t)_{\texttt{VV}}=\bar{V}_{i}(t)+\sqrt{\frac{2\log{t}}{s_{i}}},$ (3)
comprising of two distinct components: $\bar{V}_{i}(t)$, represents the estimated variance, while the second term, ${\sqrt{\frac{2\log{t}}{s_{i}(t)}}=\left(C_{t,s_{i}}\right)}$, corresponds to its associated confidence bound.
II-B Upper bound on regret of UCB-VV
First, we employ McDiarmid’s inequality to derive a concentration bound for the variance estimator $\bar{V}(n)$. Based on this, we derive an upper bound on regret of UCB-VV and show it evolves as $\mathcal{O}\log(n)$.
Lemma 1.
(Concentration Inequality for Variance Estimation): Let $X_{1},X_{2},\dots,X_{n}$ be a sequence of i.i.d. random variables bounded in $[l,u]$ with variance $\sigma^{2}$. Let $\bar{V}(n)=\frac{1}{n-1}\sum^{n}_{i=1}\left(X_{i}-\frac{1}{n}\sum_{j=1}^{n}X_{j}\right)^{2}$ be the unbiased estimator of $\sigma^{2}$. Then, the following concentration inequality holds:
$\mathbb{P}\left(\left|\bar{V}(n)-\sigma^{2}\right|>\epsilon\right)\leq 2\exp\left(\frac{-2{n}\epsilon^{2}}{(u-l)^{2}}\right).$
For random variables supported on $[0,1]$, concentration inequality for variance is presented in [35]. For the case, the domain of the rewards is $[l,u]$, the result follows immediately. Thus, the proof is omitted.
Theorem 1.
For a bandit setting with bounded rewards on $[l=0,u=1]$, the regret of UCB-VV is bounded as
$\mathcal{R}_{VV}(n)\leq 8\sum_{i:\sigma_{i}^{2}<\sigma_{*}^{2}}\frac{\log{n}}{\delta_{i}}+\left(1+\frac{\pi^{2}}{3}\right)\left(\sum_{i:\sigma_{i}^{2}<\sigma_{*}^{2}}\delta_{i}\right).$
Proof:
See Appendix VII-A. ∎
Algorithm 1 UCB-VV
Input: $S_{\rm P},K,n$
Parameter: $s_{i}(0)=0$, $\bar{X}_{i}(0)=0$, $\bar{V}_{i}(0)=0$ Output:Top 10 arms
1:  for each $t=0,1,\dots,S_{\rm P}n-1$ do
2:     Play arm $i=(t\,\text{mod}\,K)+1$.
3:     Update $s_{i}(t),\bar{X}_{i}(t)$, and $\bar{V}_{i}(t)$.
4:  end for
5:  Calculate $B_{i}(S_{\rm P}n-1)_{\texttt{VV}}$ $\forall\,i$.
6:  for each $t=S_{\rm P}n,S_{\rm P}n+1,\dots,n$ do
7:     Play arm $i=\operatorname*{argmax}\limits_{i\in\{1,2,\dots,K\}}B_{i}(t-1)_{\texttt{VV}}$.
8:     Update $s_{i}(t),\bar{X}_{i}(t)$, and $\bar{V}_{i}(t)$.
9:     Calculate $B_{i}(t)_{\texttt{VV}}$.
10:  end for
Algorithm 2 Call execution with UCB-VV
Input: Output of Algorithm 1 : Top 10 arms
Output:Rewards
1:  for each $t=0,1,\dots,N$ do
2:     for each i = 1: Top 10 do
3:        if $S_{t}>R$ then
4:           Play arm i
5:           Collect reward.
6:        else
7:           Hold
8:        end if
9:     end for
10:  end for
II-C Lower bound on regret
In this section, we establish the order optimality of the proposed algorithm UCB-VV by deriving the lower bound of the problem, and show it to be matching up to constant or logarithmic factors with the upper bound of $\mathcal{O}\log(n)$. For that, we start with a couple of basic definitions.
Definition 2.
The ${\rm KL}$ divergence between 2 distributions $\nu$ and $\tilde{\nu}$ is given by
$I(\nu,\nu^{\prime})=\mathbb{E}_{\nu}\bigg{[}\log\frac{f_{\nu}(X)}{f_{\tilde{\nu}}(X)}\bigg{]},$
where $E_{\nu}$ denotes the expectation operator with respect to $\nu$, $f_{\nu}(X)$ is the probability density function of $\nu$ and $f_{\tilde{\nu}}(X)$ is the probability density function of $\tilde{\nu}$.
Definition 3.
A policy $\pi$ is $\alpha$-consistent $10<\alpha<$ 1 if for one parameter distributions and for all $i\neq*$, $\mathbb{E}[s_{i}(n)]\leq n^{\alpha},$ where * represents the optimal arm.
Specifically, we assume that the distribution of the arm $i$ is given by $f(.;\theta_{i})$ and the distribution model $\mathcal{F}=(f(.;\theta_{1}),...,f(.;\theta_{K}))$ can be represented by ${\Theta}=(\theta_{1},...,\theta_{K})$. The parameters $\theta_{i}$ take values from a set $\mathcal{U}$ satisfying the following regularity conditions (similar to that in [36]).
Assumption 1.
For any $\theta$, $\lambda$, and $\lambda^{\prime}\in\mathcal{U}$, and for any $\epsilon>0$, there exists a $\zeta>0$ such that $0<\sigma^{2}(\lambda^{\prime})-\sigma^{2}(\lambda)<\zeta$ implies $|I(f(.;\theta),f(.;\lambda^{\prime}))-I(f(.;\theta),f(.;\lambda))|<\epsilon$. This condition ensures that small changes in the variance of distributions result in small changes in their KL-divergences. Thus, implies that the parameter space $\mathcal{U}$ is sufficiently dense to allow fine distinctions between different arms.
Since we are interested in the variance of the arms, the rewards can be transformed to have equal means without altering the relative sequence of corresponding variances. As an example, consider a scenario in which we have $2$ sequences $X$ and $Y$ with $r$ independent and identically distributed (i.i.d) random variables such that $X:X_{1},X_{2},\dots X_{r}$ and $Y:Y_{1},Y_{2},\dots Y_{r}$ with the means $\mu_{X}$ and $\mu_{Y}$, and variances $\sigma^{2}_{X}$ and $\sigma^{2}_{Y}$, respectively. From these, we construct two difference sequences $Z$ and $W$ such that $Z:Z_{1}=X_{2}-X_{1},Z_{2}=X_{3}-X_{2},\dots Z_{r-1}=X_{r}-X_{r-1}$ and $W:W_{1}=Y_{2}-Y_{1},W_{2}=Y_{3}-Y_{2},\dots W_{r-1}=Y_{r}-Y_{r-1}$. The mean of the resultant sequences is $\mu_{Z}=\mu_{W}=0$ but the variances are $\sigma^{2}_{Z}=2\sigma^{2}_{X}$ and $\sigma^{2}_{W}=2\sigma^{2}_{Y}$. We see that the variances of the transformed variables are proportional to the variance of the original variables. So, $0<\sigma^{2}_{Z}-\sigma^{2}_{W}<\zeta$ implies their distance with $f(.;\theta)$ will also be close, i.e, $|I(f(.;\theta),f(.;Z))-I(f(.;\theta),f(.;W))|<\epsilon$.
Assumption 2.
For all $\theta$ and $\lambda\in\mathcal{U}$, let $X$ be a sub-Gaussian random variable with distribution $f(.;\theta)$. The random variable $Y=f(X;\lambda)$ is sub-Gaussian. The assumption of sub-Gaussianity is required for using the Chernoff bound later in the proof, where we invoke this assumption.
We establish a preliminary lemma before proceeding to state the theorem.
Lemma 2.
Let $\Theta$ be the given distribution model, and let $i$ denote the index of a suboptimal arm under $\Theta$. Under Assumption 1, number of times arm $i$ is played under $\pi$ for any constant $C_{1}<1-\alpha$,
$\lim_{n\rightarrow\infty}\mathbb{P}_{\mathcal{F}}\bigg{[}s_{i}(n)\geq\frac{C_{1}\log n}{I(f_{i},f_{*})}\bigg{]}=1.$
Furthermore, with Assumption 2, there exists $n_{o}\in\mathbb{N}$ such that
$\mathbb{P}_{\mathcal{F}}\bigg{[}s_{i}(n)\geq\frac{C_{1}\log n}{I(f_{i},f_{*})}\bigg{]}\geq C_{2},\quad\textit{for all}\quad n>n_{0}$
where constant $10<C_{2}<$ 1 is independent of $n$ and $\mathcal{F}$.
Proof:
See Appendix VII-B. ∎
Theorem 2.
Consider the MAB problem where the objective is to optimize the variance of the observed rewards. Let $\pi$ be an $\alpha$-consistent policy and $\Theta=\left\{\left\{\theta_{i}\right\}_{i=1}^{K}\colon\theta_{i}\in\mathcal{U}\right\}$ be the distribution model. Under Assumption 1, the model-specific regret satisfies, for any constant $C_{1}<1-\alpha$, $10<C_{2}<$ 1, with $f_{i}$ and $f_{*}$ as the reward distribution of the i-th arm and the optimal arm respectively, we have
$\liminf_{n\rightarrow\infty}\frac{R(n)}{\log n}\geq\sum_{i=1,i\neq*}^{K}\frac{C_{1}C_{2}\delta_{i}}{I(f_{i},f_{*})}.$
Proof:
See Appendix VII-C. ∎
From Theorem 1 and Theorem 2, we see UCB-VV achieves an upper bound on the regret of $\mathcal{O}\log(n)$ and matches a lower bound of the same order up to constant factors, demonstrating that our proposed algorithm is asymptotically optimal.
III Problem Formulation for BAI
Next, we examine the challenge of BAI within a fixed budget. The goal of the learner is to identify and recommend the arm with the highest variance using $n$ samples. Specifically, we propose an algorithm called SHVV, that draws motivation from the foundational framework of sequential halving algorithm introduced by Karnin et al [17]. Specifically, we explore a scenario where we have a predetermined budget of $n$, aiming to maximize the likelihood of accurate identification of the highest variance arm. The total budget $n$ is divided equally into $\log_{2}(K)$ elimination rounds. Let $A_{r}$ be the set of arms in round $r$, which are pulled an equal number of times. After completing a round, we eliminate half of the arms with the lowest estimated variance. The goal of BAI setting is to minimize the error probability, i.e., recommending a sub-optimal arm at the end of the time horizon, given as,
$e_{n}=\mathbb{P}(J_{n}\neq*).$
The SHVV algorithm is given in Algorithm 3.
III-A Upper bound on error for BAI
In this section, we establish preliminary lemmas to provide the foundation for the subsequent theorem, which is the upper bound on the probability of error for SHVV, specifically focusing on the erroneous elimination of the optimal arm. The analysis hinges on inspecting the empirical third quartile of the surviving arms at each round, which serves as a critical statistic for guiding the elimination process. We first present Lemma 3, which quantifies the likelihood that any suboptimal arm $i\in A_{r}$ outperforms the optimal arm.
Lemma 3.
Assuming that the best arm is not eliminated prior to round $r$. Then for any arm $i\in A_{r}$
$\mathbb{P}\left(\bar{V}_{i}(t_{r})>\bar{V}_{*}(t_{r})\right)\leq\exp\left(-\frac{(t_{r}-1)^{2}\delta_{i}^{2}}{2t_{r}}\right)$
where $t_{r}=\frac{n}{|A_{r}|\log_{2}(K)}$.
Proof:
This probability is evaluated as
$\mathbb{P}(\bar{V}_{i}(t_{r})>\bar{V}_{*}(t_{r}))=\mathbb{P}\bigg{(}\frac{1}{t_{r}-1}\sum_{q=1}^{t_{r}}(X_{i}(q)-\bar{X}_{i})^{2}-$
$\frac{1}{t_{r}-1}\sum_{q=1}^{t_{r}}(X_{*}(q)-\bar{X}_{*})^{2}-(\sigma^{2}_{i}-\sigma^{2}_{*})>\delta_{i}\bigg{)}$
$\leq\exp\left(-\frac{(t_{r}-1)^{2}\delta_{i}^{2}}{2t_{r}}\right).$
The last inequality follows from Lemma (1). ∎
Building upon this foundation, Lemma 4 provides a complementary result by quantifying the probability that the optimal arm is erroneously eliminated during round $r$.
Lemma 4.
The probability that the best arm is eliminated in round $r$ is at most
$3\exp\left(\frac{-\left(n-4i_{r}\log_{2}(K)\right)^{2}\delta_{i_{r}}^{2}}{8ni_{r}\log_{2}(K)}\right)$
where $i_{r}=\frac{K}{2^{k+2}}$, $|A_{r}|=4i_{r}$.
Proof:
See Appendix VII-D. ∎
Leveraging the results of Lemmas 3 and 4, we now state and prove the main theorem, which establishes an upper bound on the overall probability of error.
Theorem 3.
The probability of error is upper bounded as
$e_{n}\leq 3\log_{2}(K)\exp\left(\frac{-(n-\log_{2}(K)^{2})}{8n\log_{2}(K)u^{4}H_{2}}\right)$
where $H_{2}:=max_{i\neq*}\frac{i}{\delta_{i}^{2}}$.
Proof:
Using Lemma 4 and the union bound, the best arm is eliminated in one of $\log_{2}(K)$ phases with probability at most
$3\sum_{r=1}^{\log_{2}(K)}\exp\left(-\frac{(n-4i_{r}\log_{2}(K))^{2}\delta_{i_{r}}^{2}}{8ni_{r}\log_{2}(K)}\right)$
$\leq 3\log_{2}(K)\exp\Bigg{(}-\frac{(n-4i_{r}\log_{2}(K))^{2}}{8n\log_{2}(K)}\times\frac{1}{\max_{i}i\delta_{i}^{-2}}\Bigg{)}$
$\leq 3\log_{2}(K)\exp\left(\frac{-(n-K\log_{2}(K))^{2}}{8n\log_{2}(K)H_{2}}\right).$
Algorithm 3 SHVV
Input: $K,n$
Parameter: $\bar{V}_{i}=0$ Initialize: $A_{1}\leftarrow K$, $r\leftarrow 1$
1:  for  $r=0$ to $\lceil\log_{2}K\rceil-1$  do
2:     Play arm $a_{i}\in A_{r}$ for $t_{r}=\frac{n}{|A_{r}|\log_{2}(K)}$
3:     Update $\bar{V}_{i}$.
4:     Calculate the reward
5:     Let $A_{r+1}$ be the set of $\lceil\frac{A_{r}}{2}\rceil$ arms in $A_{r}$ with largest empirical average
6:  end foroutput arm in $A_{\lceil\log_{2}K\rceil}$
III-B Lower bound on error for BAI
To establish the lower bound, we examine $K$ bandit instances, each consisting of $K$ arms but each characterized by a distinct optimal arm. The learner is aware that the bandit environment it encounters must correspond to one of these $K$ bandit settings. The lower bound is derived by proving that the learner, however good it is, will nevertheless make a mistake in identifying the bandit instance for a set of observations. Let $G^{B}$ be the setting where $B\in\{1,\dots K\}$ is optimal. Accordingly,
$G^{1};=\nu_{k}\sim U[l_{k},u_{k}]\quad\forall k\in\{1,\dots K\},$
for $0\leq l_{1}\leq l_{k}\leq u_{k}\leq u_{1}\leq 1$, i.e., the first arm has the highest variance. Let there be another bandit setting $G^{k}$:
$G^{k}$ =$=\begin{cases}\nu_{k}=\nu_{i}&i\neq k\\ \nu_{k}\sim U\left[\tilde{l}_{k},\tilde{u}_{k}\right]&i=k\end{cases}$
where $0\leq\tilde{l}_{k}\leq l_{k}\leq u_{k}\leq\tilde{u}_{k}\leq 1$. Note, ${\tilde{\sigma}^{2}_{k}}>\sigma_{1}^{2}>\sigma^{2}_{k}$.
Lemma 5.
Let $Z_{1},Z_{2},\dots,Z_{n}$ be a sequence of i.i.d. random variables such that $|Z_{s}|\leq M$ for all $s$, then for $\epsilon>0$, by Hoeffding’s inequality we have
$\mathbb{P}\left(\left|\frac{1}{t}\sum Z_{s}-\mathbb{E}[Z]\geq\epsilon\right|\right)\leq 2\exp\left(\frac{-2{n}\epsilon^{2}}{M^{2}}\right).$
Theorem 4.
For any bandit strategy that returns an arm $J_{n}$ at time $n$, for constant $M>0$, it holds that:
$\max_{1\leq k\leq K}\mathbb{P}_{k}\left(J_{n}\neq k\right)\geq$
$\hskip 56.9055pt\frac{1}{6}\exp\left(-60\frac{n}{H_{\sigma}(1)\sigma_{\nu}^{2}}-\frac{M}{\sqrt{2}}\sqrt{n\log(6nK)}\right)$
where $\mathbb{P}_{k}(J_{n}\neq k)$ is the probability of error under problem $G^{k}$, $H(k)=\sum_{1\leq i\leq K,i\neq k}(\delta_{i})^{-2}$ measures the difficulty of identifying the best arm in problem $G^{k}$, $H_{\sigma}(1)=H(1)\delta_{k}$.
Proof:
See Appendix VII-E. ∎
IV Sub-Gaussian Distributions and Sharpe Ratio
In this section, we extend our framework from bounded to sub-Gaussian distributions. In particular, we derive a concentration inequality for the sample variance of sub-Gaussian random variables, followed by extending the analysis to establish novel concentration bounds for the SR.
Theorem 5.
Let $X_{1},X_{2},\dots,X_{n}$ be i.i.d. sub-Gaussian random variables with mean $\mu=\mathbb{E}[X_{i}]$, variance $\sigma^{2}=\mathbb{E}[(X_{i}-\mu)^{2}]$, and sub-Gaussian parameter $v^{2}$, satisfying:
$\mathbb{E}[\exp(\lambda(X_{i}-\mu))]\leq\exp\left(\frac{\lambda^{2}v^{2}}{2}\right),\quad\forall\lambda\in\mathbb{R}.$
Then, for a constant $C>0$ we have:
$\mathbb{P}\left(|\bar{V}(n)-\sigma^{2}|>\epsilon\right)\leq 4\exp\left(-Cn\min\left(\frac{\epsilon^{2}}{v^{4}},\frac{\epsilon}{v^{2}}\right)\right).$
Proof:
See Appendix VII-F. ∎
Figure 1: Regret v/s time for UCB-VV, $\epsilon$-Greedy and VTS for $\delta=0.1$ and (b) $\delta=0.5$ (c) UCB-VV v/s KL-UCB.
IV-A Concentration inequality for Sharpe Ratio
The SR, defined as the ratio of expected reward to its standard deviation (or variance), is a standard metric for evaluating performance relative to risk in sequential decision-making frameworks. While our earlier research on bounded random variables derived path-dependent concentration inequalities for the SR [32], real-world scenarios often involve unbounded rewards. Extending results to the sub-Gaussian distributions not only broadens their applicability but also streamlines the analysis by removing dependencies of trajectory-specific bounds tied to bounded variables. Here, we establish a new concentration bound for the sample SR built on the variance bounds for sub-Gaussian random variables.
Theorem 6.
Let $X_{1},X_{2},\dots,X_{n}$ be i.i.d. sub-Gaussian random variables with mean $\mu=\mathbb{E}[X_{i}]$, variance $\sigma^{2}=\mathbb{E}[(X_{i}-\mu)^{2}]$, and sub-Gaussian parameter $v^{2}$. If the true SR is $S=\frac{\mu}{\sigma^{2}}$ and empirical SR is $\bar{S}(n)=\frac{\bar{\mu}(n)}{\bar{V}(n)}$, there exists a constant $c>0$ depending on $\mu,\sigma^{2},v^{2}$ such that:
$\mathbb{P}\left(|\bar{S}(n)-S|\geq\eta\right)\leq 4\exp\left(-cn\min\left(\eta^{2},\eta\right)\right).$
Proof:
See Appendix VII-G. ∎
IV-B UCB-Sharpe algorithm and its regret
[…]
Algorithm 4 UCB-Sharpe
Input: $K,n,v^{2}$
Initialize: Pull each arm once
1:  for $t=K+1$ to $T$ do
2:     for each arm $i=1,\dots,K$ do
3:        Compute sample mean $\bar{\mu}_{i}(t)$ and variance $\bar{V}_{(}t)$
4:        Compute estimated Sharpe ratio $\bar{S}_{i}(t)=\frac{\bar{\mu}_{i}}{\bar{\sigma}_{i}^{2}}$
5:        Compute confidence bonus $\beta_{i}(t)=\sqrt{\frac{1}{cs_{i}(t)}\log(4t^{2})}$
6:        Set index: $\text{Index}_{i}(t)=\bar{S}_{i}(t)+\beta_{i}(t)$
7:     end for
8:     Pull arm $I(t)=\operatorname*{argmax}_{i}\text{Index}_{i}(t)$
9:  end for
Let $S_{*}=\max_{i}S_{i}$ be the optimal SR, then the sub-optimality gap for arm $i$ is $\Delta_{i}=S_{*}-S_{i}$. The cumulative regret at time $n$ is
$\text{R}_{\rm SR}(n)=\sum_{i:\Delta_{i}>0}\Delta_{i}\cdot\mathbb{E}[s_{i}(n)],$ (4)
where $s_{i}(n)$ is the number of times arm $i$ is pulled by time $n$.
Theorem 7.
If UCB-SR is run for $K$ arms having sub-Gaussian reward distributions with parameter $v^{2}$, then its expected regret after $n$ of plays is at most
$\text{R}_{\rm SR}(n)\leq\sum_{i:\Delta_{i}>0}\left(\frac{9}{c\Delta_{i}}\log n+\mathcal{O}(1)\right).$
Proof:
See Appendix VII-H. ∎
The regret scales logarithmically with $n$, similar to classical UCB algorithms, with constants depending on the sub-Gaussian parameter $v^{2}$ and sub-optimality gap $\Delta_{i}$.
Figure 2: (a) Error probability $e_{n}$ of SHVV and uniform sampling algorithm for 6 experimental setups defined.
V Numerical results
V-A Performance for UCB-VV
In Fig. 1, we compare the performance of our proposed algorithm UCB-VV with existing algorithms like VTS, $\epsilon$-greedy and KL-UCB. We run all algorithms for $2$ arms and a time horizon of $n=5000$. In Fig. 1a, for $\delta_{i}=0.1$, we see UCB-VV outperforms $\epsilon$-greedy for $\epsilon=0.1$ as well as $\epsilon=0.2$. The regret of VTS is the lowest. In Fig. 1b, for $\delta_{i}=0.2$, we see the regret decreases for all algorithms as the distinction between optimal and sub-optimal arms gets easier. However, the drop in regret is larger for UCB-VV and VTS. UCB-VV outperforms $\epsilon$-greedy for $\epsilon=0.1$ and $\epsilon=0.2$. The regret of VTS is again the lowest among all.
In case of $\delta_{i}=0.1$, we observe the regret for $\epsilon=0.1$ is larger than for $\epsilon=0.2$. This is because a larger $\epsilon$ implies larger exploration and thus the algorithm is able to pull each arm a greater number of times and can gather more data about all the arms, before converging on the optimal arm. But excessive exploration can lead to higher regret, as we can see in the case of $\delta_{i}=0.5$. We also observe that VTS is outperforming our proposed algorithm for both sub-optimality gaps. While our proposed algorithm is optimal and offers theoretical guarantees, the superior performance of VTS highlights the need for a theoretical investigation of VTS. Thus providing the direction for future research, which could further enhance the applicability and impact of VTS. KL-UCB uses KL divergence to build confidence intervals, offering more tight bounds than traditional UCB methods that depend on simpler concentration inequalities, such as Hoeffding’s or McDiarmid’s inequality. KL-UCB is asymptotically optimal for Bernoulli bandits. This also reflects in the Fig. 1c as we compare UCB-VV with KL-UCB for $\delta=0.1$ and $\delta=0.25$ for Bernoulli distributions.
V-B Performance for SHVV
In Fig. 2, we see the performance of SHVV for a fixed budget of $n=5000$ for 6 setups. Each setup presents a unique scenario, as the sub-optimality gaps follow geometric progression and arithmetic progression or can be equal within a group of sub-optimal arms. We maintain the $\sigma_{1}^{2}=\frac{1}{12}$ as our optimal arm for all cases.
  1. 1.
    Experiment - 1: One group of sub-optimal arms: for $K=16$, $\sigma^{2}_{2:16}=\frac{1}{15}$, for $K=32$, $\sigma^{2}_{2:32}=\frac{1}{15}$, and for $K=64$, $\sigma^{2}_{2:64}=\frac{1}{15}$.
  2. 2.
    Experiment - 2: Two groups of sub-optimal arms: for $K=16$, $\left[\sigma^{2}_{2:6}=\frac{1}{14},\;\sigma^{2}_{7:16}=\frac{1}{17}\right]$, for $K=32$, $\left[\sigma^{2}_{2:14}=\frac{1}{14},\;\sigma^{2}_{15:32}=\frac{1}{17}\right]$ and for $K=64$, $\left[\sigma^{2}_{2:30}=\frac{1}{14},\;\sigma^{2}_{31:64}=\frac{1}{17}\right]$.
  3. 3.
    Experiment - 3: Arithmetic progression: $\sigma_{i}^{2}=\frac{1}{13}-0.0021(i-2)$, $i\in\{2,\dots,K\}$.
  4. 4.
    Experiment - 4: Geometric progression: $\sigma_{i}^{2}=\left(\frac{1}{12}\right)0.98^{i}$, $i\in\{2,\dots,K\}$.
  5. 5.
    Experiment - 5: Random generation: for each realization, the lower and upper limits are randomly generated with $l,\,u\in[0,1]$ such that $l<u$.
We observe that as the number of arms increases, $e_{n}$ increases. We see in Fig. 2a2b, 2c, 2d and 2f, an error corresponding to the detection of optimal arm increases as $K$ jumps from $16-32$ and $32-64$ for a fixed budget of $32e$ 3. The trend is similar for uniform sampling too, with SHVV having the lowest error. This implies that as arms grow, the budget required is greater than provided $n$, for the algorithm to perform optimally with growing arms.
Also, we see in Fig. 2e, for SHVV, as the number of arms increases, $e_{n}$ decreases as the budget allocated is also increased with the increase in arms. The error of SHVV is the least compared to uniform sampling through all experiments.
VI Case study: call option trading
Options are a widely used financial tool that investors employ to speculate on market movements or hedge against potential risks. The potential profits for option buyers are limitless, but their losses are capped, which is why they pay a premium. Conversely, option writers face potentially unlimited losses, while their profits are restricted, which is why they receive a premium. Call options are a fundamental derivative instrument in finance, granting the holder the right, but not the obligation, to purchase an underlying asset at the strike price, within a specified time period [9]. The buyer expects stock prices to rise, while the writer anticipates a decline. Volatility is a key driver of the option’s premium, as it reflects the market’s expectation of the asset’s future price fluctuations. The profit for the buyer is calculated as $\max(0,S_{t}-R)-P_{p}$, where $S_{t}$ is the stock price of the underlying asset at time $t$, $R$ is the strike price, i.e., the predetermined price at which the buyer can purchase the asset, and $P_{p}$ is the premium paid as a cost of purchasing the call option.
Figure 3: Flow Chart for call option and trade execution using SHVV and UCB-VV.
Top 8 Stocks $49$ $48$ $27$ $62$ $46$ $57$ $42$ $41$
Rewards $195.53$ $160.63$ $843.53$ $54.23$ $264.81$ $616.77$ $161.61$ $31.66$
Total Reward $2328.81$
Total Premium $622.74$
Table I: Top 8 stocks with the reward accumulated by option trading and the premium paid.
[…]
Experiment: We employ GBM to simulate stock prices, a widely adopted stochastic process in quantitative finance for its ability to model realistic stock price trajectories, making it invaluable for back-testing trading strategies, evaluating investment opportunities, and pricing financial derivatives. To ensure the analysis remains responsive to evolving market dynamics, a rolling window approach should be incorporated. This method prioritizes recent data, enabling the capture of emerging trends and volatility patterns that would otherwise be obscured or averaged out over longer time frames. By continuously updating the dataset to reflect the most current market conditions, the rolling window enhances both the relevance and accuracy of the analysis. For this study, we also used a rolling window of size $\tau=90$. This dynamic adaptation makes it especially valuable for time-sensitive financial applications, where timely insights are critical for informed decision-making.
We implement SHVV algorithm to shortlist “top 8” stocks, referring to the $8$ stocks with the highest variances given a budget of $t_{1}=20e3$. We then employ UCB-VV to select a stock for trading at time $t$ with an additional budget of $t_{2}=1e3$. If the condition $S_{t}>R$ is met, we collect the corresponding profits. This process continues iteratively: at each time step, we select the next stock according to the UCB-VV index, execute the trade if the condition holds, and accumulate the profits across all selected stocks until time horizon $n=t_{1}+t_{2}$ as shown in Table I.
We further compare the performance of our proposed framework with UCB for $n=1000$. We see in Fig 4, for $n=1000$, we achieve substantially higher cumulative rewards compared to UCB.
Figure 4: Cumulative reward (profit) v/s trading period of GBM simulated $100$ stocks for the proposed strategy v/s UCB.
VII Appendices
VII-A Proof of Theorem 1
We need to bound each $s_{i}(n)$ to bound the regret. Let $C_{t,s_{*}}=\sqrt{\frac{2\log{t}}{s_{*}}}$ be upper confidence bound. Let $z$ be any arbitrary positive integer and assume that the arm $i$ has been pulled $z$ times, including the pilot fraction. We use a subscript * to any quantity that refers to the optimal arm. At the start of the experiment, each arm is picked at least once since $S_{\rm P}=K/n$, including the sub-optimal arms. Thus,
$s_{i}(n)$ $=1+\sum_{t=K+1}^{n}\mathbb{I}\left\{\pi(t)=i\right\}$
where policy $\pi$ is UCB-VV, and $\mathbb{I}(\cdot)$ is the indicator function:
$\mathbb{I}\left\{\pi(t)=i\right\}=\begin{cases}1&\mathrm{if}\;\pi(t)=i,\\ 0&\mathrm{else.}\end{cases}$
$\text{Now},\;\;s_{i}(n)$ $\overset{(a)}{\leq}z+\sum_{t=K+1}^{n}\mathbb{I}\left\{\pi(t)=i,s_{i}\geq z\right\}$
$\overset{(b)}{\leq}z+\sum_{t=K+1}^{n}\mathbb{I}\Big{\{}\bar{V}_{*}(t-1)+C_{t-1,s_{*}}\leq$
$\hskip 56.9055pt\bar{V}_{i}(t-1)+C_{t-1,s_{i}},s_{i}\geq z\Big{\}}$
$\overset{(c)}{\leq}z+\sum_{t=K+1}^{n}\mathbb{I}\Big{\{}\min\limits_{0<s_{*}<t}\{\bar{V}_{*}(s_{*})+C_{t-1,s_{*}}\}\leq$
$\hskip 56.9055pt\max\limits_{z<s_{i}<t}\{\bar{V}_{i}(s_{i})+C_{t-1,s_{i}}\}\Big{\}}$
$\overset{(d)}{\leq}z+\sum_{t=1}^{\infty}\sum_{s_{*}=1}^{t-1}\sum_{s_{i}=z}^{t-1}\mathbb{I}\Big{\{}\bar{V}_{*}(s_{*})+C_{t,s_{*}}\leq$
$\hskip 56.9055pt\bar{V}_{i}(s_{i})+C_{t,s_{i}}\Big{\}}$ (5)
In step (a), we consider the sub-optimal arm $i$ has been picked at least $z$ times. In step (b), to pick the sub-optimal arm at time $t$, its index must be at least equal to or greater than that of the optimal arm. Step (c) involves looking back at all the optimal values we had until now, then taking their minimum, and all sub-optimal values we had, then taking their maximum. This will lead to an overcount of (b). In step (d), we upper bound by summing over all indices for the optimal and sub-optimal arms, which will include the event in step (c). Now $\bar{V}_{*}(s_{*})+C_{t,s_{*}}\leq\bar{V}_{i}(s_{i})+C_{t,s_{i}}$ implies at least one of the following must be true
$\bar{V}_{*}(s_{*})\leq\sigma^{2}_{*}-C_{t,s_{*}}$ (6)
$\bar{V}_{i}(s_{i})\geq\sigma_{i}^{2}+C_{t,s_{i}}$ (7)
$\sigma^{2}_{*}<\sigma_{i}^{2}+2C_{t,s_{i}}$ (8)
$C_{t,s_{*}}$ and $C_{t,s_{i}}$ are the bias terms for the optimal and sub-optimal arm, respectively. We bound the probability of events (6) and (7) using Lemma 1:
$\mathbb{P}\left(\bar{V}_{*}(s_{*})\leq\sigma^{2}_{*}-C_{t,s_{*}}\right)\leq e^{-2s_{*}(C_{t,s_{*}})^{2}}$
where $C_{t,s_{*}}=\sqrt{\frac{2\log{t}}{s_{*}}}$. Substitute $C_{t,s_{*}}$ above, we get
$\mathbb{P}\left(\bar{V}_{*}(r)\leq\sigma^{2}_{*}-C_{t,s_{*}}\right)$ $\leq e^{-2s_{*}\left(\sqrt{\frac{2\log{t}}{s_{*}}}\right)^{2}}=t^{-4}.$
Similarly for $C_{t,s_{i}}$, probability of (7) will be $\leq t^{-4}$. We see that the condition in equation (8) arises when the true variances of the optimal arm and sub-optimal arm are close to each other. If the true variance of optimal is less than the true variance of any sub-optimal arm with the addition of 2 times the uncertainty term, i.e., $C_{t,s_{i}(q)}$, the agent is bound to make an error until enough number of pulls have taken place. As $z$ represents the initial number of trials, and if these trials are sufficient, then equation (8) becomes false. We see that for $z=\Big{\lceil}\frac{8\log{t}}{\delta^{2}_{i}}\Big{\rceil}$, i.e., $s_{i}=z$, the number of trials of the sub-optimal arm is substantial enough to make the occurrence of an error event to become highly unlikely. This can be seen as,
$\sigma^{2}_{*}-\sigma_{i}^{2}-2C_{t,s_{i}}$ $=\sigma^{2}_{*}-\sigma_{i}^{2}-2\sqrt{\frac{2\log{t}}{s_{i}}}$
$\geq\sigma^{2}_{*}-\sigma_{i}^{2}-2\sqrt{\frac{2\log{t}\delta^{2}_{i}}{8\log{t}}}=\sigma^{2}_{*}-\sigma_{i}^{2}-\delta_{i}=0.$
It implies that in total of $n$ pulls at a certain time-step $t$, pulls of arm $i$ i.e $s_{i}\geq\frac{8\log{n}}{\delta^{2}_{i}}$ to bound the expected number of pulls for sub-optimal arm $i$ as,
$\mathbb{E}[s_{i}(n)]\leq\bigg{\lceil}\frac{8\log{n}}{\delta^{2}_{i}}\bigg{\rceil}+\sum_{t=1}^{\infty}\sum_{s_{*}=1}^{t-1}\sum_{s_{i}=\frac{8\log{n}}{\delta^{2}_{i}}}^{t-1}2t^{-4}$
$\leq\frac{8\log{n}}{\delta^{2}_{i}}+\sum_{t=1}^{\infty}\sum_{s_{*}=1}^{t}\sum_{s_{i}=1}^{t}2t^{-4}\leq\frac{8\log{n}}{\delta^{2}_{i}}+1+\frac{\pi^{2}}{3}.$
Therefore, using union bound
$\mathcal{R}_{VV}(n)\leq 8\sum_{i:\sigma_{i}^{2}<\sigma_{*}^{2}}\frac{\log{n}}{\delta_{i}}+\left(1+\frac{\pi^{2}}{3}\right)\left(\sum_{i:\sigma_{i}^{2}<\sigma_{*}^{2}}\delta_{i}\right).$
Note that for $s_{i}\geq{\frac{8\log{n}}{\delta^{2}_{i}}}$, the probability of error diminishes at a rate that contributes to logarithmic regret.
VII-B Proof of Lemma 2
Assume a distribution model $\mathcal{F}$, such that the first arm is optimal, denoted by $\ast$. We construct a new distribution model $\mathcal{F}^{i}$ by changing $f_{i}$ to $\tilde{f}_{i}$ where $i\neq\ast$ is optimal. Thus
$\mathcal{F}=(f_{1},f_{2},....f_{i}...f_{K})\quad\text{and}\quad\mathcal{F}^{i}=(f_{1},f_{2},....\tilde{f}_{i}....f_{K})$
such that $\left(\zeta>\tilde{\sigma}_{i}^{2}-\sigma^{2}_{*}>0\right)$ and $\left|I(f_{i},f_{*})-I(f_{i},\tilde{f}_{i})\right|\leq\epsilon$ for arbitrary small $\epsilon$ using assumption 1. Let $\gamma$ be the log-likelihood ratio between the $\mathcal{F}$ and $\mathcal{F}^{i}\colon\gamma=\log\frac{f_{i}\left(X_{i}\left(t_{i}(1)\right)\right)\dots f_{i}\left(X_{i}\left(t_{i}(s_{i})\right)\right)}{\tilde{f}_{i}\left(X_{i}\left(t_{i}(1)\right)\right)\dots\tilde{f}_{i}\left(X_{i}\left(t_{i}(s_{i})\right)\right)}$ where $t_{i}(s_{i})$ is the time step when the $s_{i}$-th observation of arm $i$ occurred. We show that it is unlikely to have $s_{i}<\frac{C_{1}\log n}{I(f_{i},\tilde{f}_{i})}$ under two different scenarios for $\gamma$, i.e, $\gamma>C_{3}\log n$ and $\gamma<C_{3}\log n$.
Case 1: $\gamma>C_{3}\log n$ for a constant $C_{3}>C_{1}$. We have
$\mathbb{P}_{\mathcal{F}}\bigg{[}s_{i}<\frac{C_{1}\log n}{I(f_{i},\tilde{f}_{i})},\gamma>C_{3}\log n\bigg{]}$
$=\mathbb{P}_{\mathcal{F}}\bigg{[}s_{i}<\frac{C_{1}\log n}{I(f_{i},\tilde{f}_{i})},\sum_{\tau=1}^{s_{i}}\log\frac{f_{i}(X_{k}(\tau))}{\tilde{f}_{i}(X_{k}(\tau))}>C_{3}\log n\bigg{]}$
$\leq\mathbb{P}_{\mathcal{F}}\bigg{[}\max_{t\leq\frac{C_{1}\log n}{{I(f_{i},\tilde{f}_{i})}}}\sum_{\tau=1}^{t}\log\frac{f_{i}(X_{k}(\tau))}{\tilde{f}_{i}(X_{k}(\tau))}>\frac{C_{3}\log n}{I(f_{i},\tilde{f}_{i})}I(f_{i},\tilde{f}_{i})\bigg{]}$
$\leq\mathbb{P}_{\mathcal{F}}\Bigg{[}\max_{t\leq\frac{C_{1}\log n}{{I(f_{i},\tilde{f}_{i})}}}\frac{1}{t}\sum_{\tau=1}^{t}\log\frac{f_{i}(X_{k}(\tau))}{\tilde{f}_{i}(X_{k}(\tau))}>\frac{C_{3}}{C_{1}}{I(f_{i},\tilde{f}_{i})}\bigg{]}$
By the strong law of large numbers $\frac{1}{t}\sum_{\tau=1}^{t}\log\frac{f_{i}(X_{k}(\tau))}{\tilde{f}_{i}(X_{k}(\tau))}\rightarrow I(f_{i},\tilde{f}_{i})$ a.s as $t\rightarrow\infty$. Thus, $\max_{t\leq\frac{C_{1}\log n}{{I(f_{i},\tilde{f}_{i})}}}\frac{1}{t}\sum_{\tau=1}^{t}\log\frac{f_{i}(X_{k}(\tau))}{\tilde{f}_{i}(X_{k}(\tau))}\rightarrow I(f_{i},\tilde{f}_{i})$ a.s as $n\rightarrow\infty$. So, we have
$\mathbb{P}_{\mathcal{F}}\left[\max_{t\leq\frac{C_{1}\log n}{{I(f_{i},\tilde{f}_{i})}}}\frac{1}{t}\sum_{\tau=1}^{t}\log\frac{f_{i}(X_{k}(\tau))}{\tilde{f}_{i}(X_{k}(\tau))}>\frac{C_{3}}{C_{1}}{I(f_{i},\tilde{f}_{i})}\right]\rightarrow 0$
as $n\rightarrow\infty$.
Conclusion: For $\gamma>C_{3}\log n$, by strong law of large numbers, we have:
$\mathbb{P}_{\mathcal{F}}\bigg{[}s_{i}<\frac{C_{1}\log n}{I(f_{i},\tilde{f}_{i})},\gamma>C_{3}\log n\bigg{]}\rightarrow 0,\,\text{as}\,n\rightarrow\infty$ (9)
Also, using Chernoff bound $\mathbb{P}_{\mathcal{F}}[s_{i}<\frac{C_{1}\log n}{I(f_{i},\tilde{f}_{i})},\gamma>C_{3}\log n]$ is upper bounded for finite $n$ as
$\frac{C_{1}\log n}{I(f_{i},\tilde{f}_{i})}n^{-aI(f_{i},\tilde{f}_{i})\frac{(C_{3}-C_{1})^{2}}{C_{1}}}$ (10)
Here $a$ is the Chernoff bound constant. This result, (10) is obtained from (Eq.(52) of [37]) and is given below.
$\mathbb{P}_{\mathcal{F}}\bigg{[}s_{i}<\frac{C_{1}\log n}{I(f_{i},\tilde{f}_{i})},\gamma>C_{3}\log n\bigg{]}$
$=\mathbb{P}_{\mathcal{F}}\bigg{[}s_{i}<\frac{C_{1}\log n}{I(f_{i},\tilde{f}_{i})},\sum_{s=1}^{s_{i}}\log\frac{f_{i}(X_{i}(s))}{\tilde{f}_{i}(X_{i}(s))}>C_{3}\log n\bigg{]}$
$\leq\mathbb{P}_{\mathcal{F}}\bigg{[}\max_{t<\frac{C_{1}\log n}{I(f_{i},\tilde{f}_{i})}},\sum_{s=1}^{t}\log\frac{f_{i}(X_{i}(s))}{\tilde{f}_{i}(X_{i}(s))}>C_{3}\log n\bigg{]}$
$\leq\sum_{t=1}^{\frac{C_{1}\log n}{I(f_{i},\tilde{f}_{i})}}\mathbb{P}_{\mathcal{F}}\bigg{[}\frac{1}{t}\sum_{s=1}^{t}\log\frac{f_{i}(X_{i}(s))}{\tilde{f}_{i}(X_{i}(s))}-\frac{1}{t}C_{1}\log{n}>$
$\hskip 71.13188pt\frac{1}{t}C_{3}\log{n}-\frac{1}{t}C_{1}\log{n}\bigg{]}$
$\leq\sum_{t=1}^{\frac{C_{1}\log n}{I(f_{i},\tilde{f}_{i})}}\mathbb{P}_{\mathcal{F}}\bigg{[}\frac{1}{t}\sum_{s=1}^{t}\log\frac{f_{i}(X_{i}(s))}{\tilde{f}_{i}(X_{i}(s))}-I(f_{i},\tilde{f}_{i})>$
$\hskip 71.13188pt\frac{1}{t}C_{3}\log{n}-\frac{1}{t}C_{1}\log{n}\bigg{]}$
$\leq\sum_{t=1}^{\frac{C_{1}\log n}{I(f_{i},\tilde{f}_{i})}}\exp(\frac{-a_{1}(C_{3}-C_{1})^{2})\log^{2}n}{t})$ (11)
$\leq\frac{C_{1}\log n}{I(f_{i},\tilde{f}_{i})}n^{-aI(f_{i},\tilde{f}_{i})\frac{(C_{3}-C_{1})^{2}}{C_{1}}}$ (12)
(11) holds because $I(f_{i},\tilde{f}_{i})>\frac{1}{t}C_{1}\log{n}$ and (12) holds according to Chernoff bound.
Case 2: $\gamma\leq C_{3}\log n$. By Markov’s Inequality, we have
$\mathbb{P}_{\mathcal{F}^{i}}\left[s_{i}<\frac{C_{1}\log n}{I(f_{i},\tilde{f}_{i})}\right]$ $=\mathbb{P}_{\mathcal{F}^{i}}\left[n-s_{i}\geq n-\frac{C_{1}\log n}{I(f_{i},\tilde{f}_{i})}\right]$
$\leq\frac{\mathbb{E}_{\mathcal{F}^{i}}[n-s_{i}]}{n-\frac{C_{1}\log n}{I(f_{i},\tilde{f}_{i})}}$ (13)
[…]
where $S(n)$ is the set of all the observations over $n$ that satisfy a particular event and $\gamma$ is the log-likelihood ratio between $\mathcal{F}$ and $\mathcal{F}^{i}$. Finally, (14) is an application of Radon-Nikodym theorem [38]. Using (13) and (14),
$\mathbb{P}_{\mathcal{F}}\left[s_{i}<\frac{C_{1}\log n}{I(f_{i},\tilde{f}_{i})},\gamma\leq C_{3}\log n\right]$
$=\mathbb{E}_{\mathcal{F}}\left[\mathbb{I}\left\{s_{i}<\frac{C_{1}\log n}{I\left(f_{i},\tilde{f}_{i}\right)},\gamma\leq C_{3}\log n\right\}\right]$
$=\mathbb{E}_{\mathcal{F}^{i}}\left[\mathbb{I}\left\{s_{i}<\frac{C_{1}\log n}{I\left(f_{i},\tilde{f}_{i}\right)},\gamma\leq C_{3}\log n\right\}\times e^{\gamma}\right]$
$\overset{(a)}{\leq}n^{C_{3}}\mathbb{E}_{\mathcal{F}^{i}}\left[\mathbb{I}\left\{s_{i}<\frac{C_{1}\log n}{I(f_{i},\tilde{f}_{i})},\gamma\leq C_{3}\log n\right\}\right]$
$\leq n^{C_{3}}\mathbb{P}_{\mathcal{F}^{i}}\left[{s_{i}<\frac{C_{1}\log n}{I\left(f_{i},\tilde{f}_{i}\right)}}\right]\leq\frac{n^{C_{3}}\mathbb{E}_{\mathcal{F}^{i}}\left[n-s_{i}\right]}{n-\frac{C_{1}\log n}{I(f_{i},\tilde{f}_{i})}}$
$\leq\frac{Kn^{C_{3}+\alpha}}{{n-\frac{C_{1}\log n}{I(f_{i},\tilde{f}_{i})}}}$ (15)
Step (a) follows from substituting $\gamma\leq C_{3}\log{n}$ in $e^{\gamma}$. The last inequality is due to $\alpha$-consistency. From (9), (15) and the fact that ${I(f_{i},f_{*})}-{I(f_{i},\tilde{f}_{i})}$ can be arbitrarily small, for $C_{3}<1-\alpha$, we have,
$\mathbb{P}_{\mathcal{F}}\bigg{[}s_{i}<\frac{C_{1}\log n}{I(f_{i},f_{*})}\bigg{]}\rightarrow 0\,\text{as}\,n\rightarrow\infty$
[…]
Using (10) and (15), we conclude, for $C_{3}<1-\alpha$, when Assumption 2 is satisfied,
$\mathbb{P}_{\mathcal{F}}\bigg{[}s_{i}<\frac{C_{1}\log n}{I(f_{i},f_{*})}\bigg{]}$ $\leq\frac{C_{1}\log n}{I(f_{i},\tilde{f}_{i})}n^{-aI(f_{i},\tilde{f}_{i})\frac{(C_{3}-C_{1})^{2}}{C_{1}}}+$
$\hskip 56.9055pt\frac{Kn^{C_{3}+\alpha}}{{n-\frac{C_{1}\log n}{I(f_{i},\tilde{f}_{i})}}}.$
Thus, there is a $n_{0}$ such that for $n>n_{0}$, $\mathbb{P}_{\mathcal{F}}\bigg{[}s_{i}\geq\frac{C_{1}\log n}{I(f_{i},f_{*})}\bigg{]}\geq C_{2}$, for some constant $C_{2}>0$ independent of $n$ and $\mathcal{F}$. We emphasize that constant $C_{1}$ and $C_{3}$ are chosen to satisfy $C_{1}<C_{3}<1-\alpha$.
VII-C Proof of Theorem 2
We need to establish a lower bound on expected sub-optimal pulls, i.e, $\mathbb{E}[s_{i}]$, which is a straightforward consequence of Lemma 2. By using Markov’s inequality, we have
$\mathbb{E}[s_{i}]\geq\mathbb{P}\left[s_{i}\geq\frac{C_{1}\log n}{I(f_{i},f_{*})}\right]\frac{C_{1}\log n}{I(f_{i},f_{*})}$
[…]
$=\frac{C_{1}}{I(f_{i},f_{*})}$
Similarly, by Markov’s inequality, we have, when assumption 2 is satisfied, there is $n_{0}\in n$ such that for all $n>n_{0}$,
$\mathbb{E}[s_{i}]\geq\frac{C_{1}C_{2}\log n}{I(f_{i},f_{*})}$
[…]
$\implies\;\liminf_{n\rightarrow\infty}\frac{R(n)}{\log n}$ $\geq\sum_{i=1,i\neq*}^{K}\frac{C_{1}C_{2}\delta_{i}}{I(f_{i},f_{*})}$
VII-D Proof of Lemma 4
$A_{r}^{\prime}$ is the set of arms in $A_{r}$, excluding the $\frac{1}{4}|A_{r}|=\frac{K}{2^{k+2}}$ arms with the largest variance and letting $D_{r}$ denote the number of arms in $A_{r}^{\prime}$ whose empirical variance is larger than that of the optimal arm,
$\mathbb{E}[D_{r}]$ $=\sum_{i\in A_{r}^{\prime}}\mathbb{P}(\bar{V}_{1}(t_{r})<\bar{V}_{i}(t_{r}))\leq\sum_{i\in A_{r}^{\prime}}\exp\left(-\frac{(t_{r}-1)^{2}\delta_{i}^{2}}{2t_{r}}\right)$
$\leq|A_{r}^{\prime}|\max_{i\in A_{r}^{\prime}}\exp\left(-\frac{\left(\frac{n}{|A_{r}|\log_{2}(K)}-1\right)^{2}\delta_{i}^{2}}{\frac{2n}{|A_{r}|\log_{2}(K)}}\right)$
$=|A_{r}^{\prime}|\max_{i\in A_{r}^{\prime}}\exp\left(-\frac{({n-|A_{r}|\log_{2}(K))^{2}}\delta_{i}^{2}}{2n|A_{r}|\log_{2}(K)}\right)$
$\leq|A_{r}^{\prime}|\exp\left(-\frac{({n-4i_{r}\log_{2}(K))^{2}}\delta_{i_{r}}^{2}}{8ni_{r}\log_{2}(K)}\right)$
Using Markov’s Inequality, we can bound the probability that $D_{r}$ exceeds a fraction of $|A_{r}^{\prime}|$. Markov’s Inequality states that for any non-negative random variable $X$ and constant $a>0$, $\mathbb{P}(X\geq a)\leq\frac{\mathbb{E}[X]}{a}.$ Therefore,
$\mathbb{P}\left(D_{r}>\frac{1}{3}|A_{r}^{\prime}|\right)\leq 3\frac{\mathbb{E}[D_{r}]}{|A_{r}^{\prime}|}\leq 3\mathbb{E}[D_{r}]$
$=3\exp\left(\frac{-\left(n-4i_{r}\log_{2}(K)\right)^{2}\delta_{i_{r}}^{2}}{8ni_{r}\log_{2}(K)}\right).$
VII-E Proof of Theorem 4
Step 1: Concentration of empirical ${\rm KL}$ divergence
A bounded distribution is defined on the interval $[l,u]$. The ${\rm KL}$ divergence between 2 such distributions $\nu$ and $\tilde{\nu}$ depends on their PDFs as given below:
[…]
Let $X_{i}(s)$ denote the samples drawn from arm $i$, which corresponds to arm in the first bandit setting $G^{1}$ The empirical ${\rm KL}$-divergence is given as:
$\hat{I}_{i}(t)=\frac{1}{t}\sum_{s=1}^{t}\log\left(\frac{f_{\nu_{i}}(X_{i}(s))}{f_{\tilde{\nu}_{i}}(X_{i}(s))}\right)$
To ensure the empirical ${\rm KL}$ divergence $\hat{I}_{i}(t)$ concentrates around true ${I}_{i}$, we define a high-probability event $\xi$:
$\xi\!\!=\!\!\left\{\forall 1\leq i\leq K,1\leq t\leq n,\big{|}\hat{I}_{i}(t)-{I}_{i}\big{|}\!\leq\!\frac{M}{\sqrt{2}}\sqrt{\frac{\log(6nK)}{t}}\right\}.$
This means that for all arms $i$ and all times $t$, the empirical ${\rm KL}$ divergence $\hat{I}_{i}(t)$ is close to the true ${\rm KL}$ divergence ${I}_{i}$, with a deviation bounded by $\frac{M}{\sqrt{2}}\sqrt{\frac{\log(6nK)}{t}}$. It ensures that the empirical observations collected by the learner are reliable enough to distinguish between the true distributions. We prove a lemma that $\mathbb{P}_{k}(\xi)$ is a high-probability event.
Lemma 6.
It holds that $\mathbb{P}_{k}(\xi)\geq\frac{5}{6}$.
To derive the above, we need to upper bound the log-likelihood ratio $\log\left(\frac{f_{\nu_{i}}(x)}{f_{\tilde{\nu}_{i}}(x)}\right)$. ${f_{\nu_{i}}(x)}$ and $f_{\tilde{\nu}_{i}}(x)$ are PDFs of bounded distributions, we assume, the log-likelihood ratio is also bounded by some constant $M>0$, depending on the specific forms of $f_{\nu_{i}}(x)$ and $f_{\tilde{\nu}_{i}}(x)$. Let $M=\sup_{x\in[l,u]}\left|\log\frac{f_{\nu_{i}}(x)}{f_{\tilde{\nu}_{i}}(x)}\right|$. Using lemma 5, for $Z_{s}=\log\frac{f_{\nu_{i}(x)}}{f_{\tilde{\nu}_{i}(x)}},\mathbb{E}[Z_{s}]={I}_{i}$, we can re-write Hoeffding’s inequality as $\mathbb{P}\left(\left|\hat{I}_{i}(t)-{I}_{i}\geq\epsilon\right|\right)\leq 2\exp\left(\frac{-2{n}\epsilon^{2}}{M^{2}}\right)$. We ensure that the KL divergence is also bounded by M, from the event $\xi$, defined above.
Let $\epsilon=C_{4}\sqrt{\frac{\log(6nK)}{n}}$, where $C_{4}$ depends on $M$,
$\mathbb{P}\left(\left|\hat{I}_{i}(t)-{I}_{i}\right|\geq C_{4}\sqrt{\frac{\log(6nK)}{n}}\right)\leq$
$\hskip 85.35826pt2\exp\left(\frac{-2{t}C_{4}^{2}\frac{\log(6nK)}{t}}{M^{2}}\right)$
$=2\exp\left(\frac{-2C_{4}^{2}\log(6nK)}{M^{2}}\right)$
Choose $C$ such that $\frac{2C_{4}^{2}}{M^{2}}=1$. Thus, $C_{4}=\frac{M}{\sqrt{2}}$. Therefore,
$\mathbb{P}\left(\left|\hat{I}_{i}(t)-{I}_{i}\right|\geq\frac{M}{\sqrt{2}}\sqrt{\frac{\log(6nK)}{t}}\right)\leq\frac{1}{6nK}$
A union bound over all arms and over all time, we get
$\mathbb{P}\left(\exists i,t:\left|\hat{I}_{i}(t)-{I}_{i}\right|\geq\frac{M}{\sqrt{2}}\sqrt{\frac{\log(6nK)}{t}}\right)\leq\frac{Kn}{6nK}=\frac{1}{6}.$
The probability that the event holds is $\mathbb{P}_{k}({\xi})\geq 1-\frac{1}{6}=\frac{5}{6}$.
Step 2: Change of Measure
Let $(n_{i})_{1\leq i\leq K}$ denote the numbers of samples collected by the agent on each arm of the bandits. These quantities are stochastic, but it holds that $\sum_{1\leq i\leq K}n_{i}=n$ by the definition of the fixed budget setting. Let us write for any $0\leq i\leq K$, $t_{i}=\mathbb{E}_{1}n_{i}$, where $\mathbb{E}_{1}(.)$ denotes the expectation under problem $G^{1}$. It also holds that $\sum_{1\leq i\leq K}t_{i}=n$. Using the change of measure identity, we relate the probabilities of error under different bandit problems. From [15], we have, for any measurable event $E$ and for any $2\leq k\leq K$:
$\mathbb{P}_{k}(E)=\mathbb{E}_{1}\left[\mathbb{I}\{E\}\exp\left(-n_{k}\cdot\hat{{I}}_{k}(n_{k})\right)\right],$ (16)
Consider an event $E_{k}=\{J_{n}=1\}\cap\{\xi\}\cap\{n_{k}\leq 6t_{k}\}$, where:
  • $J_{n}=1$: The learner outputs arm $1$ as the best arm,
  • $\xi$: A high-probability event where empirical ${\rm KL}$ divergences concentrate,
  • $n_{k}\leq 6t_{k}$: Arm $k$ is pulled at most $6t_{k}$ times.
From equation 16, we have
$\mathbb{P}_{k}(E_{k})=\mathbb{E}_{1}\left[\mathbb{I}\{E\}\exp\left(-n_{k}\cdot\hat{{I}}_{k}(n_{k})\right)\right]$
$\geq\exp\left[-6t_{k}{I}_{k}-\frac{M}{\sqrt{2}}\sqrt{n\log(6nK)}\right]\mathbb{P}_{1}[E_{k}]$ (17)
Step 3: Lower Bound on $\mathbb{P}_{1}[E_{k}]$.
Assume that for the algorithm that we consider
$\mathbb{E}_{1}(J_{n}\neq 1)\leq\frac{1}{2}$ (18)
i.e., that the probability that the algorithm makes a mistake on problem $1$ is less than $\frac{1}{2}$. For any $2\leq i\leq K$, it holds by Markov’s inequality that
$\mathbb{P}_{1}(n_{i}\geq 6t_{i})\leq\frac{\mathbb{E}_{1}n_{i}}{6t_{i}}=\frac{1}{6}$ (19)
Using (18) and (19), Lemma 5 ($\mathbb{P}_{k}(\xi)\geq\frac{5}{6}$) and using union bounds for any $2\leq k\leq K$,
$\mathbb{P}_{1}(E_{k})\geq 1-\bigg{(}\frac{1}{6}+\frac{1}{2}+\frac{1}{6}\bigg{)}=\frac{1}{6}.$
Using (17) and the fact that for any $2\leq k\leq K,\mathbb{P}_{k}(J_{n}\neq k)\geq\mathbb{P}_{k}(E_{k})$ implies that for any $2\leq k\leq K$
$\mathbb{P}_{k}\left({J_{n}\neq{k}}\right)\geq\frac{1}{6}\exp\left[-6t_{k}{I}_{k}-\frac{M}{\sqrt{2}}\sqrt{n\log(6nK)}\right]$
KL-divergence in case of variance for uniform distribution is:
$I_{k}=\frac{1}{2}\log\frac{f_{\nu}(x)}{f_{\tilde{\nu}}(x)}=\frac{1}{2}\log{\frac{\sigma_{\tilde{\nu}}^{2}}{\sigma_{\nu}^{2}}}\leq\frac{1}{2}\left(\frac{\sigma_{\tilde{\nu}}^{2}}{\sigma_{\nu}^{2}}-1\right)=\frac{\delta_{k}}{2\sigma_{\nu}^{2}}$ (20)
Step 4: Final bound
Since $\sum_{2\leq i\leq K}\delta_{i}^{-2}=H(1)$ and $\sum_{1\leq i\leq K}n_{i}=n$, then there exists $2\leq k\leq K$ such that $t_{k}=\frac{n}{H(1)\delta_{k}^{2}}$. Combining the results from Steps 1–3, we derive the lower bounds on the error probability in terms of $H(1)$, which is given as;
$\mathbb{P}_{k}\left({J_{n}\neq{k}}\right)$ $\geq\frac{1}{6}\exp\left[-6\frac{n}{H(1)\delta_{k}^{2}}\frac{\delta_{k}}{2\sigma_{\nu}^{2}}-\frac{M}{\sqrt{2}}\sqrt{n\log(6nK)}\right]$
$=\frac{1}{6}\exp\left[-3\frac{n}{H_{\sigma}(1)\sigma_{\nu}^{2}}-\frac{M}{\sqrt{2}}\sqrt{n\log(6nK)}\right].$
Note: For Bernoulli distribution, lower bound proves order-optimality and can be followed from [16] with some minor changes in bounded support of $\nu_{k}$ and $\tilde{\nu}_{k}$.
VII-F Proof of Theorem 5
Proof:
Rewriting $\bar{V}(n)$ in terms of the true mean $\mu$:
$\bar{V}(n)={\frac{1}{n}\sum_{i=1}^{n}(X_{i}-\mu)^{2}}-{(\bar{\mu}(n)-\mu)^{2}}.$
Applying the triangle inequality, we have
$|\bar{V}(n)-\sigma^{2}|\leq\left|\frac{1}{n}\sum_{i=1}^{n}(X_{i}-\mu)^{2}-\sigma^{2}\right|+(\bar{\mu}(n)-\mu)^{2}.$
Let $Y_{i}=(X_{i}-\mu)^{2}$. Since $X_{i}$ is sub-Gaussian with parameter $v^{2}$, $Y_{i}$ is sub-exponential with parameters $(v^{2},4v^{2})$. Applying Bernstein’s inequality for sub-exponential random variables,
$\mathbb{P}\left(\left|\frac{1}{n}\sum_{i=1}^{n}Y_{i}-\sigma^{2}\right|>\epsilon\right)\leq 2\exp\left(-C_{5}n\min\left(\frac{\epsilon^{2}}{v^{4}},\frac{\epsilon}{v^{2}}\right)\right),$
for some constant $C_{5}>0$. For the second term, since $\bar{\mu}(n)$ is sub-Gaussian with parameter $v^{2}/n$. Therefore,
$\mathbb{P}\left(|\bar{\mu}(n)-\mu|>\epsilon_{1}\right)\leq 2\exp\left(-\frac{n\epsilon_{1}^{2}}{2v^{2}}\right).$
Setting $\epsilon_{1}=\sqrt{\epsilon/2}$, we get $\mathbb{P}\left((\bar{\mu}(n)-\mu)^{2}>\frac{\epsilon}{2}\right)\leq 2\exp\left(-\frac{n\epsilon}{4v^{2}}\right).$ Combining both terms:
$\mathbb{P}\left(|\bar{V}(n)-\sigma^{2}|>\epsilon\right)\leq 2\exp\left(-C_{5}n\min\left(\frac{\epsilon^{2}}{v^{4}},\frac{\epsilon}{v^{2}}\right)\right)+$
$2\exp\left(-\frac{n\epsilon}{4v^{2}}\right).$
Thus, $\exists C:0<C<\min\left\{\frac{1}{4},C_{5}\right\}$ such that the statement of the theorem follows. ∎
VII-G Proof of Theorem 6
Proof:
We begin by expressing the deviation of the empirical SR from its true value:
$|\bar{S}(n)-S|=\left|\frac{\bar{\mu}(n)}{\bar{V}(n)}-\frac{\mu}{\sigma^{2}}\right|=\left|\frac{\bar{\mu}(n)\sigma^{2}-\mu\bar{V}(n)}{\bar{V}(n)\sigma^{2}}\right|.$
By the triangle inequality, we decompose the equation as,
$|\bar{S}(n)-S|\leq\frac{|\bar{\mu}(n)-\mu|}{\bar{V}(n)}+\frac{|\mu|\cdot|\bar{V}(n)-\sigma^{2}|}{\bar{V}(n)\sigma^{2}}.$
We first leverage concentration inequalities to stabilize the denominator by enforcing its constancy, i.e., to show with high probability that $\bar{V}(n)\geq\frac{\sigma^{2}}{2}$. Utilizing the result of Theorem 5, where we put $\epsilon=\frac{\sigma^{2}}{2}$ and get:
$\mathbb{P}\left(\bar{V}(n)\geq\frac{\sigma^{2}}{2}\right)\leq 2\exp\left(-Cn\min\left(\frac{\sigma^{4}}{4v^{4}},\frac{\sigma^{2}}{2v^{2}}\right)\right).$
Thus with probability at least $1-2\exp\left(-Cn\min\left(\frac{\sigma^{4}}{4v^{4}},\frac{\sigma^{2}}{2v^{2}}\right)\right)$, we have $\bar{V}(n)\geq\frac{\sigma^{2}}{2}\,\Rightarrow\,\frac{1}{\bar{V}(n)}\leq\frac{2}{\sigma^{2}}$. For this event, the deviation bound becomes:
$|\bar{S}(n)-S|\leq\frac{2|\bar{\mu}(n)-\mu|}{\sigma^{2}}+\frac{2|\mu|\cdot|\bar{V}(n)-\sigma^{2}|}{\sigma^{4}}.$
To ensure $|\bar{S}(n)-S|\leq\eta$, it suffices to require both terms on the right-hand side to be at most $\eta/2$. That is,
$|\bar{\mu}(n)-\mu|\leq\frac{\eta\sigma^{2}}{2},\quad|\bar{V}(n)-\sigma^{2}|\leq\frac{\eta\sigma^{4}}{2|\mu|}.$
Using the concentration bound for the sample mean $\mathbb{P}\left(|\bar{\mu}(n)-\mu|\geq\epsilon\right)\leq 2\exp\left(-\frac{n\epsilon^{2}}{2v^{2}}\right)$, we set $\epsilon=\frac{\eta\sigma^{2}}{2}$ to obtain:
$\mathbb{P}\left(|\bar{\mu}(n)-\mu|\geq\frac{\eta\sigma^{2}}{2}\right)\leq 2\exp\left(-\frac{n\eta^{2}\sigma^{4}}{8v^{2}}\right).$
Similarly, invoking the sub-exponential tail bound stated in Theorem 5 for the deviation $|\bar{V}(n)-\sigma^{2}|$ and setting $\epsilon_{1}=\frac{\eta\sigma^{4}}{4|\mu|}$, we obtain
$\mathbb{P}\left(|\bar{V}(n)-\sigma^{2}|\geq\frac{\eta\sigma^{4}}{2|\mu|}\right)\leq$
$\hskip 56.9055pt2\exp\left(-C_{6}n\min\left(\frac{\eta^{2}\sigma^{8}}{4\mu^{2}v^{4}},\frac{\eta\sigma^{4}}{2|\mu|v^{2}}\right)\right),$
for some constant $C_{6}>0$. Combining these bounds via the union bound, the total probability of deviation satisfies:
$\mathbb{P}\left(|\bar{S}(n)-S|\geq\eta\right)$ $\leq 2\exp\left(-\frac{n\eta^{2}\sigma^{4}}{8v^{2}}\right)+$
$2\exp\left(-C_{6}n\min\left(\frac{\eta^{2}\sigma^{8}}{4\mu^{2}v^{4}},\frac{\eta\sigma^{4}}{2|\mu|v^{2}}\right)\right).$
Each exponential term decays at a rate governed by $n\min(\eta^{2},\eta)$, up to constants depending on $\mu,\sigma^{2},v^{2}$. Absorbing all constants into a single parameter $c>0$, we conclude the statement of the theorem. ∎
VII-H Proof of Theorem 6
Proof:
The proof largely follows the corresponding concentration inequality and is in line with other proofs of UCB-type algorithms. Hence, here we provide an abridged sketch of the proof. From the concentration inequality on sample SR, the estimated SR $\bar{S}_{i}(t)$ satisfies:
$\mathbb{P}\left(|\bar{S}_{i}(t)-S_{i}|>\beta_{i}(t)\right)\leq\frac{1}{t^{2}},$
Applying a union bound over all arms $i$ up to $n$:
$\mathbb{P}\left(\exists i,t\leq n:|\bar{S}_{i}(t)-S_{i}|>\beta_{i}(t)\right)\leq\sum_{i=1}^{K}\sum_{t=1}^{n}\frac{1}{t^{2}}\leq\frac{\pi^{2}}{6}K.$
Thus, with high probability, for all $i$ and $t$, we have:
$S_{i}\leq\bar{S}_{i}(t)+\beta_{i}(t),\quad\bar{S}_{i}(t)\leq S_{i}+\beta_{i}(t).$
Suppose a suboptimal arm $i$ (with $\Delta_{i}>0$) is selected at time $t$. By the UCB-Sharpe algorithm, its index must satisfy:
$\bar{S}_{i}(t)+\beta_{i}(t)\geq\bar{S}_{i_{*}}(t)+\beta_{i_{*}}(t)$
$\Rightarrow S_{i}+2\beta_{i}(t)\geq S_{*}-\beta_{i_{*}}(t)\;\Rightarrow\Delta_{i}\leq 3\beta_{i}(t)$
Solve for $s_{i}(t)$ using $\beta_{i}(t)=\sqrt{\frac{1}{cs_{i}(t)}\log(4t^{2})}$, we get
$\Delta_{i}\leq 3\sqrt{\frac{1}{cs_{i}(t)}\log(4n^{2})}\quad\Rightarrow\quad s_{i}(t)\leq\frac{9}{c\Delta_{i}^{2}}\log(4n^{2}).$
Summing over all suboptimal arms:
$\mathbb{E}[s_{i}(n)]\leq\frac{9}{c\Delta_{i}^{2}}\log(4n^{2})+\mathcal{O}(1).$
Therefore, using (4), we get the statement of the theorem. ∎